OpenStudy (anonymous):

Im trying to write this expression as a single fraction 2+ (a-8b/2a)-(3a-3b/8a)

OpenStudy (anonymous):

\[2+ {a-8b\over2a}-{3a-3b\over8a}\]find a common denominator \((8a)\):\[{16a\over8a}+{4a-32b\over8a}-{3a-3b\over8a}\]put them all under one denominator:\[{(16a)+(4a-32b)-(3a-3b)\over8a}\]distribute the positive and negative signs:\[{16a+4a-32b-3a+3b\over8a}\]combine the like terms in the numeraotr:\[17a-29b\over8a\]

OpenStudy (campbell_st):

well find the common denominator... it should be 8a so you have \[\frac{2(8a)}{8a} + \frac{4(a - 8b)}{8a} - \frac{3a - 3b)}{8a}\] now you can simplify the numerators///

OpenStudy (anonymous):

hope that's helpful :) @learning1